6) You are trying to decide from three investment opportunities with the following end-of-year CFs. Which investment would you choose assuming a 10% discount rate? Remember to show what each investment is worth to you today, to explain your choice. End of Year 400 400 400 400 500 400 400 400 400 600 6 600
Answer #1

The option which gives the maximum net present value will be chosen

Present value factors

= 1 / (1 + r) ^ n

Where,

r = Rate of interest = 10% or 0.10

n = Years = 0 to 7

So, PV Factor for year 2

= 1 / (1.10^2)

= 1 / 1.21

= 0.826446

The following table shows the calculations:

Option A
Calculations Years 0 1 2 3 4 5 6 7
A Cash Flow 0 400 400 400 400 0 0 0
B PV Factors 1 0.909091 0.826446 0.751315 0.683013 0.620921 0.564474 0.513158
C = A x B Present Values 0 363.6364 330.5785 300.5259 273.2054 0 0 0
D = Sum C Net Present value 1267.95
Option B
Calculations Years 0 1 2 3 4 5 6 7
A Cash Flow 0 0 0 0 400 400 400 400
B PV Factors 1 0.909091 0.826446 0.751315 0.683013 0.620921 0.564474 0.513158
C = A x B Present Values 0 0 0 0 273.2054 248.3685 225.7896 205.2632
D = Sum C Net Present value 952.63
Option C
Calculations Years 0 1 2 3 4 5 6 7
A Cash Flow 0 500 0 0 600 0 0 600
B PV Factors 1 0.909091 0.826446 0.751315 0.683013 0.620921 0.564474 0.513158
C = A x B Present Values 0 454.5455 0 0 409.8081 0 0 307.8949
D = Sum C Net Present value 1172.25

So, as per above calculations, Investment A should be chosen as it gives the maximum net present value